Calculate the value of g at latitude 60° N.
What is its value at the poles ?
Values of physical constants :
g = 10 ms-2
G = 6.7 x 10-11 Nm2 lcg-2
Mass of the Sun = 2 x 1030 kg
Re = 6.4 x 106 m
1
Expert's answer
2018-06-21T08:55:08-0400
Overall variation of g with latitude:
g (θ) = 9.78031846 (1+ 0.0053024 sin ² θ – 0.0000058 sin² 2θ)
Thus, the value of g at latitude 60° N:
g (60°)= 9.78031846 (1+ 0.0053024 sin^2⁡〖60°〗 – 0.0000058 sin^2⁡〖2(60°)〗 )=9.819 m/s^2
Thus, the value of g at the poles:
g (90°)= 9.78031846 (1+ 0.0053024 sin^2⁡〖90°〗 – 0.0000058 sin^2⁡〖2(90°)〗 )=9.832 m/s^2